From the 1/14th Daily Journal Summaries for the month of April, 1968.

April 15-21, 1968

15 April 1968


GOLDEN DRAGONS The morning hours found Co's B and D moving east noon, B and D Co's were lifted from PZ's to assmue security missions on LZ 14 and LZ Mile Highe respectively. Both units were OPCON to 1/35 Inf. In the late evening Co C detected movement on the southwest edge of the perimeter and thre a hand grenade. Negative results were reported.

16 April 1968


GOLDEN DRAGONS continued operations around LZ No Slack. RCN platoon moved to the firebase in order to be at an LZ for movement to LZ Swinger. Airstrikes were placed on LZ Swinger (LZ Construction).

Plans: Multiple combat assault will be conducted on LZ Swinger after additional airstrikes.


17 April 1968


GOLDEN DRAGONS conducted a combat assault with Charlie company followed by Alpha Company and the Hvy Mort Platoon. Delta reverted OPCON 1/14 in from 1/35 at 1630.

PLANS: Continued effort to clear the fire base of the heavy bamboo and prepare positions for the artillery.

18 April 1968


Bde Intel report:

1st suspected regiment, 1700m radius, loc YA 830921, time: early afternoon

2nd unidentified, 1200m radius, loc YA 837935, time: late afternoon

3rd suspected regiment, 3700m radius, loc YA 845933, time: late afternoon


GOLDEN DRAGONS continued construction of it's new fire base, LZ Swinger. Charlie Company found 140 - 82 MM mortar rounds dropped along a trail approximately 200 meters below the firebase. Bravo Company reverted OPCON of the GOLDEN DRAGONS. Hqs(-) closed LZ Swinger (except TOC) from Polei Kleng.

PLANS: Continue construction and sweep out from the firebase with Bravo to the West, Alpha to the East and Delta to the South.

19 April 1968


GOLDEN DRAGONS continued to improve their new firebase, LZ Swinger. All elements in the AO conducted local patrols around their positions to increase security for the fire base. At approximately 2035 hours "B" Co detected movement in front of their perimeter. Arty was called with neg. results.

PLANS: Continue to improve LZ SWINGER. Alpha Co to move Southwest to sweep the valley south of LZ SWINGER. Other elements will continue to improve their positions and conduct local patrols.

20 April 1968


GOLDEN DRAGONS continued work on LZ SWINGER. A Co conducted a sweep of the valley east of LZ SWINGER with negative results. D Co pushed south toward the southern valley with negative results.

PLANS: Continue to work area around LZ SWINGER with B Co to the west, A Co to the east and D Co to the south.

21 April 1968


GOLDEN DRAGONS continued operations around LZ SWINGER. Alpha Co completed it's sweep of the western valley finding only one filter for an NVA gas mask. Delta Co continued to move south toward the low ground south of LZ SWINGER. At 2019 hrs Brave Co received one incoming round. Believed to be 122mm rocket. Negative casualties.

PLANS: Continue operations in assigned AO. All elements to improve their defensive positions and continue local patrols.

APRIL 22-30